OpenStudy (anonymous):

please i need help with this Transportation The gas tank in Tony’s car holds 15 gallons, and the car can travel 25 miles for each gallon of gas. When Tony begins with a full tank of gas, the function f (x) = -__1 25 x + 15 gives the amount of gas f(x) that will be left in the tank after traveling x miles (if he does not buy more gas). Graph this function and give its domain and range.

OpenStudy (anonymous):

Function=4. Domain=5. Range=3. Please click the medal next to my name.

OpenStudy (anonymous):

Next to my answer that is.

OpenStudy (anonymous):

x must be nonnegative, since you can't travel a negative number of miles. y must also be nonnegative, since you can't have a negative amount of gas. Tony starts with 15 gallons when x=0, and this amount will decrease all the way down to zero, so the range is 0 <= f(x) <= 15. f(x) = 0 when -1/25x + 15 = 0, or 15 = x/25, so x = 15*25 = 375. Therefore the domain is 0 <= x <= 375.